# Title keyword 5 for citation Title for citation Title for citation

### v2

@@ 6.1.2 Intelligent Industrial Robots
Industrial robots will be widely utilized in future industrial manufacturing and transportation scenarios [287]. For instance, construction robots can complete building tasks under remote control. Industrial robots will be considered to replace human workers for hazardous tasks, particularly in extreme working environments such as high temperatures and high altitudes. This will enable an ”unmanned factory,” where people only need to use their terminal devices. After completing the AI training of intelligent robots, it is essential to validate the AI models before deployment to ensure the safe and efficient execution of production tasks by the intelligent robots.
6G AIaaS provides AI training services for robots, such as data collection through robots, applying model training for multi-agent learning, and distributing the trained models to the robots. The extreme transmission capabilities of 6G will support rapid model training and parameter exchange between robots and the network. 6G AIaaS also offers multi-system linkage capabilities, allowing various terminal devices to network with robots flexibly and enabling remote control via computers, smartphones, and VR devices. The coverage of the 6G mobile communication system extends from traditional user terminals to environmental sensing devices. Robots can connect and interact with the environment through 6G AIaaS’s sensing technology. They can even engage in cross-environment and cross-system learning with robots from different departments, sharing experiences. Additionally, the 6G network can leverage DT networks to provide AI verification services, iteratively training AI models with greater robustness and better performance through performance pre-validation. 6G AIaaS offers AI verification services for trained models, requiring the DT network to generate sample data in more scenarios than the physical environment, reducing the overhead and performance impact of data collection from the physical network, thereby further improving the stability and performance of AI models.
@@ 6.1.3 Precision Medical
Healthcare is closely related to individual life and health safety, making the accuracy and reliability of predictive diagnostics crucial. Traditional predictive diagnostics based on single AI models face limitations in accuracy, reliability, and real-time performance due to the maturity of AI algorithms and the availability of case data samples. By leveraging AI over 6G networks, geographically dispersed medical institutions and individual practitioners can connect more extensively, aggregate more AI models and case data samples, and rapidly transmit and synchronize information between doctors and patients. This enhances FL and group learning, continually improving the accuracy, reliability, and real-time performance of predictive diagnostics and treatment actions, thereby increasing trust in the prospects of smart healthcare.
Smart healthcare can also achieve more rational planning and optimized allocation of various medical resources through 6G network AI, significantly reducing the physical and psychological stress on doctors and patients and addressing issues such as ”difficulties in seeking medical treatment,” ”long waiting times,” and ”over-medication.” Additionally, protecting the privacy of predictive diagnostics and treatment results for certain diseases is vital. Using a 6G network, AI-distributed technology can better safeguard individual medical information privacy.
@@ 6.1.4 Ultra-Intelligent Transportation
In the future, ultra-intelligent transportation will involve various stages across the temporal dimension, including high-definition map downloads, vehicle environment perception, environmental prediction, and route planning [289]. Spatially, ultra-intelligent transportation will cover scenarios such as highways, urban roads, and enclosed areas. The vision for ultra-intelligent transportation is to enable intelligent vehicles to provide users with a safer, more efficient, and more comfortable riding experience at a lower cost. As the new foundational infrastructure for future information and communication technology, the 6G system boasts ultra-low latency, high-reliability transmission, supercomputing power, and precise sensing. These characteristics will support the real-time transmission, processing, and intelligent decision-making of data related to ultra-intelligent transportation, achieving full network intelligence.
@@
Targeting sea, land, air, and space transportation, the 6G network’s extensive coverage and intelligent capabilities enable autonomous driving and smart traffic management for vehicles, intelligent scheduling for ships and aviation, and unified control and management of water, land, underground, and aerial transportation. With the 6G network’s deep learning and multi-sensor fusion capabilities, an integrated perception and decision-making architecture is established for vehicles, roadside infrastructure, and cloud platforms. This architecture advances a new era of intelligent transportation characterized by intelligent travel, ubiquitous services, and comprehensive management and control. In terms of travel, it achieves the organic unification of individual optimal travel solutions and the overall highest efficiency of the transportation system.
@@ 6.2 QoAIS
Service Category Indicator Dimension QoAIS Indicators
AI Services Performance Boundary of performance metrics, training time, generalization, reusability, robustness, interpretability, consistency between loss function and optimization objectives, fairness, etc.
@@
QoAIS refers to a system of metrics framework for AI services in the network, which broadly needs to involve dimensions such as performance, connectivity, computation, data, security, and orchestration [290]. The performance dimension metrics in QoAIS differ from the model KPIs for ML. This is because the latter refers to the evaluation metrics of the model itself, which are mainly used for optimization and tuning of the model. The former extends that content, such as generalizability and robustness of the model/algorithm, to describe the whole system more comprehensively in terms of performance. The metrics of other dimensions in QoAIS are new metrics that consider the 6G network itself and thus extend it. As shown in Table 7, the content of the whole QoAIS system should include:
@@ 7 Standardization Process
As wireless communication and AI integration is a focus of the industry’s research direction, important standardization organizations and industry alliances at home and abroad have launched research on integrating mobile communication networks and AI since 2017. These include the third-generation Mobile Communication Partnership Project (3GPP), International Telecommunication Union (ITU), European Telecommunication Standardization Institute (ETSI), China Communication Standardization Association (CCSA), IMT-2020 (5G) Promotion Group, IMT-2030 (6G) promotion group, etc. Currently, 3GPP has three Technical Specification Groups: the core network and terminals (CT), radio access network (RAN), and Service and system aspects (SA). Each TSG has 4 to 6 Working Groups. The 3GPP SA WG2 (responsible for developing the overall 3GPP system architecture and services, including User Equipment, Access Network, Core Network and IP multimedia Subsystem System Architecture and Services) and the 3GPP RAN WG3 (responsible for the overall radio network architecture and the specification of protocols for the related network interfaces) already accomplished some specification work on data collection, data analytic, and associated procedures at core network and radio network. 3GPP SA WG5 has progressed in standardizing network management data analytics services for managing AI/ML Models and functions. ETSI’s zero-touch network and service management working group focuses on the research and analysis of technologies related to automated closed-loop operations, aiming to achieve a high degree of automation in network and service management. ETSI’s Experiential Networked Intelligence (ENI) working group is focused on defining an experiential and aware network management architecture that uses AI to improve operators’ experience in network deployment and operations. Its goal is to use AI technology to improve the efficiency and automation of network management by defining a perception-adaptive-decision-execute control model. Figure 27 displays the application process of the AI model.
@@
3GPP has conducted relevant research on network intelligence and automation and introduced network data analytics function (NWDAF) as a core network AI element since Release-15. NWDAF can collect data from any network functions in 5GC and third-party application functions. Due to time limitations, only slice-specific network data analytics is supported in Rel-15. Various network data and third-party application data from AF and 5GC network functions could be used as input to NWDAF, and output analytic results from NWDAF could be used by other NFs to guide network optimization, such as the following use cases been introduced in the Release-16: customized mobile management, 5G quality of service (QoS) enhancement, user plane function (UPF) selection, network functions load balancing, slicing SLA guarantee, etc. Rel-17 allows multiple NWDAFs to be deployed in a distributed architecture. Another enhancement is decomposing the monolithic tWDAF into a Model Training Logic Function, Analytics Logical Function, and Analytics Data Repository Function. Further, the data collection coordination function (DCCF) enhances the data collection framework. Rel-18, the model performance monitoring, multi-vendor ML Model sharing, Horizontal FL, and six more use cases have been standardized.
@@
In wireless access networks, 3GPP has identified typical application scenarios for SON and has applied machine learning techniques, successfully verifying its usefulness in optimizing network performance. In 2018, 3GPP SA5 Group set up a research project, ”a study on SON for 5G,” aimed at the integration of SON and 5G network technology, studying to promote the integration of SON and 5G network infrastructure by strengthening the collection and application of wireless big data.
@@
The RAN1 Working Group identified the Study Item ”Study on AI/machine learning for NR air interface” in the Rel-18 phase. The SI explores how AI/ML algorithms can enhance air interface performance (throughput, robustness, accuracy, or reliability) and reduce overhead. The main application scenarios of this research include AI-based CSI feedback, AI-based beam management, AI-based positioning enhancement, and life cycle management of AI/ML model/functionality for the air interface.
@@
The 3GPP RAN2 Working Group identified the rStudy on AI/ML for mobility in NR” research project in December 2023. The study project aims to evaluate potential benefits and potential specification impacts of AI/ML-aided mobility for network-triggered L3-based handover, considering the following aspects: AI/ML-based RRM measurement and event prediction for L3 Mobility, HO failure/RLF prediction, and Measurement events prediction.
@@
The 3GPP RAN3 Working Group identified the study item ”Study on further enhancement for data collection” in July 2020; the SI aims to enable an intelligent functional architecture for RAN through use cases to enhance data collection further and identify potential standardization implications for current NG-RAN nodes and interfaces. In February 2022, the SI was completed, and the intelligent typical functional architecture on the RAN side, input, output, feedback, and potential standard impact of three high-priority use cases (e.g., network energy saving, load balancing, and mobility optimization) were written into TR 37.817. The intelligent universal functional architecture on the RAN side includes functions such as data collection, model training, model inference, and action execution. The 3GPP RAN3 Working Group identified the ”AI/Machine Learning for NG-RAN” work item. The WI aims to incorporate data collection enhancements and signaling support specified in existing NG-RAN interfaces and architectures to enable AI/ML-based network energy saving, load balancing, and mobility optimization.
@@
The TSG SA (responsible for the overall architecture and service capabilities of 3GPP systems and the cross-3GPP TSG coordination) identified the Study item, “3GPP AI/ML Consistency Alignment,” in June 2024. The SI will investigate ongoing AI/ML work in TSG CT, TSG RAN and TSG SA Working Groups and identify instances of any potential misalignment and/or inconsistencies. TSG SA leads the study in close collaboration and inputs from TSG CT and TSG RAN. The technical report of this SI will be finalized by June 2025.
@@
In addition to the 3GPP-defined NWDAF network element, the 5G-Monarch project in Europe introduced RAN-DAF, an independent AI analysis function explicitly designed for CU in 5G NR, for data analysis and decision-making. As a network element for AI and data analysis on the RAN side, RAN-DAF can monitor and collect UE and RAN data, and AI can prioritize local processing based on these real-time data to solve the need for rapid response of operations such as wireless resource management. MoNArch recommends that the RAN-DAF pass information to the controller on the RAN side to collaborate with the RCA to optimize wireless side performance, such as slice-sensitive selection, live wireless resource control, cross-slice resource management, and so on.
@@
ITU-T Study Group 13 established the ITU-T Focus Group on Machine Learning for Future Networks and 5G in November 2017. The group has drafted several technical reports and specifications for machine learning in future networks, including network architectures, interfaces, protocols, algorithms, etc. There are three working groups (WG) under the focus group where WG1 studies potential machine learning use cases in future networks and identifies the needs of use cases, and WG2 mainly works to classify machine learning in mobile communication networks and define the required data formats and related mechanisms to protect security and privacy. WG3 primarily studies the requirements of machine learning on network architecture in mobile communication networks, including network functions, interfaces, resources, etc.
@@
In recent years, standardization organizations have also focused on the research and preparation of 6G. In June 2019, the China IMT-2030 (6G) promotion group was established to gather the strength of Chinese industry, universities and research institutes, promoting China’s 6G mobile communication technology research and carrying out international exchanges and cooperation and successively releasing several white papers and research reports such as the data format and model proposal of mobile communication and AI integration. In June 2023, the 44th meeting of the ITU radio communication sector 5D working group (ITU-R WP5D) was held in Geneva, Switzerland, and the recommendation on the framework and overall objectives of the IMT for 2030 and Beyond was completed. This proposal aims to set a framework and overall goals for the development of international mobile communications (IMT) in 2030 and beyond, involving a global 6G vision consensus, 6G goals and trends, typical 6G scenarios and capability indicators system, including the convergence of AI and communication scenarios, and define AI-related indicators. On September 12, 2024, at the 105th Plenary meeting held in Melbourne, Australia, 3GPP officially launched the first 6G Study Item - “6G use cases and Service Requirements”. This initiative formally marks the 6G technology from the pre-research stage into the substantive standardization stage.

### v3

@@ 5.1.2 Intelligent Industrial Robots
Industrial robots will be widely utilized in future industrial manufacturing and transportation scenarios [253]. For instance, construction robots can complete building tasks under remote control. Industrial robots will be considered to replace human workers for hazardous tasks, particularly in extreme working environments such as high temperatures and high altitudes. This will enable an “unmanned factory”, where people only need to use their terminal devices. After completing the AI training of intelligent robots, it is essential to validate the AI models before deployment to ensure the safe and efficient execution of production tasks by the intelligent robots.
6G AIaaS provides AI training services for robots, such as data collection through robots, applying model training for multi-agent learning, and distributing the trained models to the robots. The extreme transmission capabilities of 6G will support rapid model training and parameter exchange between robots and the network. 6G AIaaS also offers multi-system linkage capabilities, allowing various terminal devices to network with robots flexibly and enabling remote control via computers, smartphones, and VR devices. The coverage of the 6G mobile communication system extends from conventional user terminals to environmental sensing devices. Robots can connect and interact with the environment through 6G AIaaS’s sensing technology. They can even engage in cross-environment and cross-system learning with robots from different departments, sharing experiences. Additionally, the 6G network can leverage DT networks to provide AI verification services, iteratively training AI models with greater robustness and better performance through performance pre-validation. 6G AIaaS offers AI verification services for trained models, requiring the DT network to generate sample data in more scenarios than the physical environment, reducing the overhead and performance impact of data collection from the physical network, thereby further improving the stability and performance of AI models.
@@ 5.1.3 Precision Medical
Healthcare is closely related to individual life and health safety, making the accuracy and reliability of predictive diagnostics crucial. conventional predictive diagnostics based on single AI models face limitations in accuracy, reliability, and real-time performance due to the maturity of AI algorithms and the availability of case data samples. By leveraging AI over 6G networks, geographically dispersed medical institutions and individual practitioners can connect more extensively, aggregate more AI models and case data samples, and rapidly transmit and synchronize information between doctors and patients. This enhances FL and group learning, continually improving the accuracy, reliability, and real-time performance of predictive diagnostics and treatment actions, thereby increasing trust in the prospects of smart healthcare.
Smart healthcare can also achieve more rational planning and optimized allocation of various medical resources through 6G network AI, significantly reducing the physical and psychological stress on doctors and patients and addressing issues such as “difficulties in seeking medical treatment,” “long waiting times”, and “over-medication”. Additionally, protecting the privacy of predictive diagnostics and treatment results for certain diseases is vital. Using a 6G network, AI-distributed technology can better safeguard individual medical information privacy.
@@ 5.1.4 Ultra-Intelligent Transportation
In the future, ultra-intelligent transportation will involve various stages across the temporal dimension, including high-definition map downloads, vehicle environment perception, environmental prediction, and route planning. Spatially, ultra-intelligent transportation will cover scenarios such as highways, urban roads, and enclosed areas. The vision for ultra-intelligent transportation is to enable intelligent vehicles to provide users with a safer, more efficient, and more comfortable riding experience at a lower cost. As the new foundational infrastructure for future ICT, the 6G system boasts ultra-low latency, high-reliability transmission, supercomputing power, and precise sensing. These characteristics will support the real-time transmission, processing, and intelligent decision-making of data related to ultra-intelligent transportation, achieving full network intelligence.
@@
Targeting sea, land, air, and space transportation, the 6G network’s extensive coverage and intelligent capabilities enable autonomous driving and smart traffic management for vehicles, intelligent scheduling for ships and aviation, and unified control and management of water, land, underground, and aerial transportation. With the 6G network’s DL and multi-sensor fusion capabilities, an integrated perception and decision-making architecture is established for vehicles, roadside infrastructure, and cloud platforms. This architecture advances a new era of intelligent transportation characterized by intelligent travel, ubiquitous services, and comprehensive management and control. In terms of travel, it achieves the organic unification of individual optimal travel solutions and the overall highest efficiency of the transportation system.
@@ 5.2 QoAIS
QoAIS refers to a system of metrics framework for AI services in the network, which broadly needs to involve dimensions such as performance, connectivity, computation, data, security, and orchestration [255]. The performance dimension metrics in QoAIS differ from the model KPI for ML. This is because the latter refers to the evaluation metrics of the model itself, which are mainly used for optimization and tuning of the model. The former extends that content, such as generalizability and robustness of the model/algorithm, to describe the whole system more comprehensively in terms of performance. The metrics of other dimensions in QoAIS are new metrics that consider the 6G network itself and thus extend it. As shown in Table 3, the content of the whole QoAIS system should include:
@@
Indicator Dimension QoAIS Indicators
Performance Boundary of performance metrics, training time, generalization, reusability, robustness, interpretability, consistency between loss function and optimization objectives, fairness, etc.
@@
5.3 Key Capabilities for Supporting AIaaS
The realization of AIaaS essentially means that the entire network serves as a platform capable of providing AI services for users and network devices, which requires the support of many key capabilities. This section will enumerate these key capabilities and potential technologies from the following three aspects.
1) Interconnection of User-Data-Model: To construct an AIaaS platform, it is essential to have extensive data support. This can be achieved by leveraging the network’s inherent sensing capabilities, along with the data collection abilities of numerous user terminals and IoT devices, to enable real-time data acquisition and storage. Furthermore, while ensuring data privacy, it is crucial to establish a direct connection between users and data. This approach provides essential data resources for model training and inference services within the AIaaS platform. Secondly, AI services rely on AI models, and different services require the design of appropriate algorithms to implement intelligent data processing and decision-making processes. Users can utilize the model resources provided by the network or customize models as needed. Therefore, a dedicated interface should be established to connect users with models, providing necessary support for customizing AI models.
2) Management of Heterogeneous Resources and Task Scheduling: 6G networks feature ubiquitous distributed computing power, with CPU/GPU capabilities and cache resources spread across various CN nodes, BS, user devices, and third-party equipment. Using software defined network, computing resources at different nodes can be discovered and tagged, and resources can be uniformly measured through virtualization. Further employing distributed resource management protocols, such as distributed hash tables, enables efficient storage and retrieval of computing resource information. This approach also facilitates automatic updates and timely processing of resource information. Under the premise of resource visibility, the rational allocation of computing, caching, and bandwidth resources significantly impacts the performance of AI services. Techniques such as DRL algorithms, model predictive control, game theory, and dynamic caching can be employed to achieve adaptive resource allocation and service scheduling based on priority or fairness. These methods ensure optimal service performance across varying service loads.
3) Security Assurance in the AIaaS Lifecycle: With the ubiquitous distribution of cloud-edge computing, the demand for interconnecting computing nodes has intensified, complicating the security landscape of network connections. The dynamic scheduling of computing power introduces new requirements for point-to-point network connections and dynamic connection mechanisms across systems, domains, and even borders. To ensure the security of AI services throughout their lifecycle, comprehensive protection of data security, computational security, model security, and privacy is essential. To address these challenges, the constructed AIaaS platform should implement secure data acquisition protocols and device authentication mechanisms to ensure the legitimacy and integrity of data sources. This involves employing secure data fusion algorithms to validate and cleanse data from various terminal devices. Techniques such as homomorphic encryption can be utilized to perform aggregation operations without compromising data privacy. For large-scale data storage, effective storage and timely processing are essential. Techniques like differential privacy and FL should be employed to defend against model inversion and membership inference attacks, thereby safeguarding model parameters and training data security. Additionally, in AI services, the trained models possess significant value and intellectual property. The AIaaS platform should implement mechanisms to protect these models from unauthorized access and misuse. This includes dynamic protection strategies, such as monitoring data flow and model outputs, to safeguard against illegal activities.
In the practical application of AIaaS platforms, requirements such as data management, computational scheduling, and security protection are not isolated; they must be considered across multiple dimensions. These platforms should provide a multidimensional service experience tailored to user-specific needs. The advent of 6G networks will transcend convetional pipeline limitations by offering AIaaS services through an inherently intelligent network architecture. This will address the digital and intelligent needs of the network itself, various industries, and consumers, truly enabling the era of intelligent inclusivity.
@@ 6 Standardization Process
As wireless communication and AI integration is a focus of the industry’s research direction, important standardization organizations and industry alliances at home and abroad have launched research on integrating mobile communication networks and AI since 2017. These include the 3GPP, ITU, european telecommunication standardization institute (ETSI), china communication standardization association, IMT-2020 (5G) promotion group, IMT-2030 (6G) promotion group, etc. Currently, 3GPP has three technical specification groups (TSG): the CN, RAN, and service and system aspects (SA). Each TSG has 4 to 6 WG. The 3GPP SA WG2 (responsible for developing the overall 3GPP system architecture and services, including user equipment, access network, CT and IP multimedia subsystem system architecture and services) and the 3GPP RAN WG3 (responsible for the overall radio network architecture and the specification of protocols for the related network interfaces) already accomplished some specification work on data collection, data analytic, and associated procedures at CN and radio network. 3GPP SA WG5 has progressed in standardizing network management data analytics services for managing AI/ML Models and functions. ETSI’s zero-touch network and service management working group focuses on the research and analysis of technologies related to automated closed-loop operations, aiming to achieve a high degree of automation in network and service management. ETSI’s experiential networked intelligence (ENI) working group is focused on defining an experiential and aware network management architecture that uses AI to improve operators’ experience in network deployment and operations. Its goal is to use AI technology to improve the efficiency and automation of network management by defining a perception-adaptive-decision-execute control model. Figure 26 displays the application process of the AI model.
@@
3GPP has conducted relevant research on network intelligence and automation and introduced NWDAF as a CN AI element since Release-15. NWDAF can collect data from any network functions in 5GC and third-party application functions. Due to time limitations, only slice-specific network data analytics is supported in Rel-15. Various network data and third-party application data from AF and 5GC network functions could be used as input to NWDAF, and output analytic results from NWDAF could be used by other NFs to guide network optimization, such as the following use cases been introduced in the Release-16: customized mobile management, 5G QoS enhancement, UPF selection, network functions load balancing, slicing SLA guarantee, etc. Rel-17 allows multiple NWDAFs to be deployed in a distributed architecture. Another enhancement is decomposing the monolithic tWDAF into a Model Training Logic Function, Analytics Logical Function, and Analytics Data Repository Function. Further, the data collection coordination function enhances the data collection framework. Rel-18, the model performance monitoring, multi-vendor ML Model sharing, Horizontal FL, and six more use cases have been standardized.
@@
In wireless access networks, 3GPP has identified typical application scenarios for SON and has applied ML techniques, successfully verifying its usefulness in optimizing network performance. In 2018, 3GPP SA5 Group set up a research project, “a study on SON for 5G,” aimed at the integration of SON and 5G network technology, studying to promote the integration of SON and 5G network infrastructure by strengthening the collection and application of wireless big data.
@@
The RAN1 Working Group identified the study item “Study on AI/ML for NR air interface” in the Rel-18 phase [256]. The SI explores how AI/ML algorithms can enhance air interface performance (throughput, robustness, accuracy, or reliability) and reduce overhead. The main application scenarios of this research include AI-based CSI feedback, AI-based beam management, AI-based positioning enhancement, and life cycle management of AI/ML model/functionality for the air interface.
@@
The 3GPP RAN2 Working Group identified the “Study on AI/ML for mobility in NR” research project in December 2023 [257]. The study project aims to evaluate potential benefits and potential specification impacts of AI/ML-aided mobility for network-triggered L3-based hand off, considering the following aspects: AI/ML-based RRM measurement and event prediction for L3 Mobility, HO failure/RLF prediction, and Measurement events prediction.
@@
The 3GPP RAN3 Working Group identified the study item “Study on further enhancement for data collection” in July 2020 [258]; the SI aims to enable an intelligent functional architecture for RAN through use cases to enhance data collection further and identify potential standardization implications for current NG-RAN nodes and interfaces. In February 2022, the SI was completed, and the intelligent typical functional architecture on the RAN side, input, output, feedback, and potential standard impact of three high-priority use cases (e.g., network energy saving, load balancing, and mobility optimization) were written into TR 37.817. The intelligent universal functional architecture on the RAN side includes functions such as data collection, model training, model inference, and action execution. The 3GPP RAN3 Working Group identified the “AI/ML for NG-RAN” work item [259]. The WI aims to incorporate data collection enhancements and signaling support specified in existing NG-RAN interfaces and architectures to enable AI/ML-based network energy saving, load balancing, and mobility optimization.
@@
The TSG SA (responsible for the overall architecture and service capabilities of 3GPP systems and the cross-3GPP TSG coordination) identified the Study item, “3GPP AI/ML Consistency Alignment,” in June 2024 [260]. The SI will investigate ongoing AI/ML work in TSG CT, TSG RAN and TSG SA working groups and identify instances of any potential misalignment and/or inconsistencies. TSG SA leads the study in close collaboration and inputs from TSG CT and TSG RAN. The technical report of this SI will be finalized by June 2025.
@@
In addition to the 3GPP-defined NWDAF network element, the 5G-Monarch project in Europe introduced RAN-DAF, an independent AI analysis function explicitly designed for control unit plane in 5G NR, for data analysis and decision-making. As a network element for AI and data analysis on the RAN side, RAN-DAF can monitor and collect UE and RAN data, and AI can prioritize local processing based on these real-time data to solve the need for rapid response of operations such as wireless resource management. MoNArch recommends that the RAN-DAF pass information to the controller on the RAN side to collaborate with the RCA to optimize wireless side performance, such as slice-sensitive selection, live wireless resource control, cross-slice resource management, and so on.
@@
ITU-T Study Group 13 established the ITU-T Focus Group on ML for Future Networks and 5G in November 2017. The group has drafted several technical reports and specifications for ML in future networks, including network architectures, interfaces, protocols, algorithms, etc. There are three WG under the focus group where WG1 studies potential ML use cases in future networks and identifies the needs of use cases, and WG2 mainly works to classify ML in mobile communication networks and define the required data formats and related mechanisms to protect security and privacy. WG3 primarily studies the requirements of ML on network architecture in mobile communication networks, including network functions, interfaces, resources, etc.
@@
In recent years, standardization organizations have also focused on the research and preparation of 6G. In June 2019, the China IMT-2030 (6G) promotion group was established to gather the strength of Chinese industry, universities and research institutes, promoting China’s 6G mobile communication technology research and carrying out international exchanges and cooperation and successively releasing several white papers and research reports such as the data format and model proposal of mobile communication and AI integration. In June 2023, the 44th meeting of the ITU-R working group 5D was held in Geneva, Switzerland, and the recommendation on the framework and overall objectives of the IMT for 2030 and Beyond was completed. This proposal aims to set a framework and overall goals for the development of IMT in 2030 and beyond, involving a global 6G vision consensus, 6G goals and trends, typical 6G scenarios and capability indicators system, including the convergence of AI and communication scenarios, and define AI-related indicators. On September 12, 2024, at the 105th Plenary meeting held in Melbourne, Australia, 3GPP officially launched the first 6G Study Item - “6G use cases and Service Requirements”. This initiative formally marks the 6G technology from the pre-research stage into the substantive standardization stage.
Figure 27: The standardization process of AI/ML for wireless networks